Please explain to me how to multiply 0.5 by 0.2. Thanks for your help!
Found 2 solutions by dkppathak, stanbon:
Answer by dkppathak(439)

.5x.2=.10
or we can say answer =0.1

Answer by stanbon(75887)

Please explain to me how to multiply 0.5 by 0.2. Thanks for your help!
----
(5/10)(2/10) = 10/100 = 1/10
----
0.5*0.2 = (1/2)(0.2) = 0.1 = 1/10
----
0.5*0.2 = 0.10 = 0.1 = 1/10
==================================
Cheers,
Stan H.
===============